OpenStudy (anonymous):

The number of fish (f) in a fish tank increases when the number of algae blooms (a) decreases. Write the correct equation for this scenario and solve for the number of fish when there are 4 algae blooms. Algae Blooms | Fish 10 10 25 4

jimthompson5910 (jim_thompson5910):

slope: m = (y2 - y1)/(x2 - x1) m = (4 - 10)/(25 - 10) m = -6/15 m = -2/5

jimthompson5910 (jim_thompson5910):

hopefully you can see how I got the slope

jimthompson5910 (jim_thompson5910):

if not, then ask about any step if so, then I'll move on to the next part